### v1

# Stochastic Gradient Descent Jittering for Inverse Problems: Alleviating the Accuracy-Robustness Tradeoff

Abstract
Inverse problems aim to reconstruct unseen data from corrupted or perturbed measurements. While most work focuses on improving reconstruction quality, generalization accuracy and robustness are equally important, especially for safety-critical applications. Model-based architectures (MBAs), such as loop unrolling methods, are considered more interpretable and achieve better reconstructions. Empirical evidence suggests that MBAs are more robust to perturbations than black-box solvers, but the accuracy-robustness tradeoff in MBAs remains underexplored. In this work, we propose a simple yet effective training scheme for MBAs, called SGD jittering, which injects noise iteration-wise during reconstruction. We theoretically demonstrate that SGD jittering not only generalizes better than the standard mean squared error training but is also more robust to average-case attacks. We validate SGD jittering using denoising toy examples, seismic deconvolution, and single-coil MRI reconstruction. The proposed method achieves cleaner reconstructions for out-of-distribution data and demonstrates enhanced robustness to adversarial attacks.
Introduction
Inverse problems (IPs) aim to recover the underlying signal ${\bm{x}}\in\mathbb{R}^{n}$ from corrupted observation ${\bm{y}}\in\mathbb{R}^{m}$, where $m$ is often less than $n$, leading to non-unique solutions. IPs are typically modeled as:
${\bm{y}}=\bm{A}{\bm{x}}+{\bm{z}},$ (1)
where $\bm{A}\in\mathbb{R}^{m\times n}$ is a known forward model for linear IPs, and ${\bm{z}}$ represents unknown noise in the observation.
Machine learning (ML) algorithms show significantly better reconstructions than the classical optimization methods. The majority of the work in the field aims to promote reconstruction quality using different deep-learning models, such as transformers (Fabian, Tinaz, and Soltanolkotabi 2022) and diffusion models (Whang et al. 2022). In particular, model-based architectures (MBAs) such as loop unrolling (LU) networks inspired by classical optimization methods, unroll the objective function into a sequence of trainable iterative steps with the forward model $\bm{A}$ involved in every iteration. MBAs are considered to be more interpretable than neural networks that learn a direct inverse without using $\bm{A}$ (Monga, Li, and Eldar 2021), and achieve state-of-the-art reconstructions in different fields, such as natural image restoration (Gilton, Ongie, and Willett 2021) and medical image reconstruction (Bian 2024).
While reconstruction quality is crucial, robustness and generalization accuracy are equally important for IPs. Robustness refers to the model’s ability to handle unwanted noise, while generalization accuracy measures how well the model performs on unseen data. For example, in medical image reconstruction, a model must generalize well to detect subtle details like small bone fractures and be robust enough to eliminate artifacts caused by measurement noise. Despite their importance, robustness and generalization have not been thoroughly explored in the context of IPs, particularly in MBAs. While some studies provide empirical assessments of robustness (Gilton, Ongie, and Willett 2021; Genzel, Macdonald, and März 2022), there is a noticeable lack of theoretical analysis (Gottschling et al. 2020).
In this work, we model MBA training as a bilevel optimization problem without making specific assumptions about neural network architectures. We theoretically analyze its generalization and robustness risks in solving denoising problems and demonstrate that standard mean-squared error (MSE) training may perform poorly in both metrics.
Technologies like adversarial training (AT) (Fawzi, Moosavi-Dezfooli, and Frossard 2016; Ford et al. 2019; Kannan, Kurakin, and Goodfellow 2018) and input noise injection (Fawzi, Moosavi-Dezfooli, and Frossard 2016; Ford et al. 2019; Kannan, Kurakin, and Goodfellow 2018) are commonly used to enhance robustness, while approaches such as weight perturbation (Wu, Xia, and Wang 2020), stochastic training (Hardt, Recht, and Singer 2016) and sharpness-aware minimization (Foret et al. 2020) are designed to improve generalization. However, there is often a tradeoff between robustness and accuracy (Tsipras et al. 2018; Zhang et al. 2023; Gottschling et al. 2020; Krainovic, Soltanolkotabi, and Heckel 2023). Although strategies like careful loss function design (Pang et al. 2022; Kannan, Kurakin, and Goodfellow 2018; Zhang et al. 2019; Zheng, Zhang, and Mao 2021), data augmentation (Zhang et al. 2017), and robust architectural choices such as dropout (Yang et al. 2020) and teacher-student models (Zhao, Wang, and Wei 2024) have been explored to mitigate this tradeoff in classification tasks, the challenge of balancing accuracy and robustness in IP solvers still remains an open question (Gottschling et al. 2020). Recent work by Krainovic, Soltanolkotabi, and Heckel (2023) introduces jittering in ${\bm{y}}$ for black-box IP neural networks, achieving a comparable level of worst-case robustness to AT but with significant accuracy loss.
To bridge this gap, this work aims to improve robustness through a novel training scheme for MBAs, while maintaining generalization accuracy for IPs. Note that while specially designed architectures can also improve robustness or generalization, they are not discussed here as this work focuses on investigating different training schemes.
Our Contributions
In this work, we propose a stochastic gradient descent (SGD) jittering scheme in training MBAs. This method unrolls the objective function for solving IPs using a gradient descent algorithm, with small, random noises added to the gradient updates at each iteration to improve the training process. During inference, these jittering noises are removed, allowing the model to perform as it would in standard evaluation settings. This technique has demonstrated greater robustness and better generalization than traditional MSE training, while also being more time-efficient than AT. Our contribution can be summarized as follows,
  • (network assumptions) We investigate the performance of MBAs with general non-convex neural networks, where training often results in non-unique parameters. This analysis is crucial for demonstrating improvements across various metrics, particularly in robustness and generalization accuracy.
  • (robustness) We theoretically prove that SGD jittering implicitly promotes average-case robustness than regular MSE training. In the experiments, we further demonstrate the effectiveness of SGD jittering under worst-case attacks.
  • (generalization) We are the first to analyze generalization accuracy for IPs with small perturbations in the testing data, rather than relying on PAC-Bayes Theorem, which hypothetically promotes generalization through Hessian regularization.
  • (convergence) We also demonstrate that under the assumption of Lipschitz continuity, SGD jittering converges, ensuring the training is feasible.
Related Work
Analysis of Robustness
Smoothness is closely related to the robustness of neural networks. Works like Salman et al. (2019); Cohen, Rosenfeld, and Kolter (2019); Lecuyer et al. (2019); Li et al. (2018) show that regularization in smoothness promotes the robustness of classifiers. On the other hand, the analysis of IP is very different from classifications. Krainovic, Soltanolkotabi, and Heckel (2023) analyzes the worst-case robustness for IPs using a simple linear reconstruction model and directly solves for robustness risk. However, optimizing a convex model results in a solution that excels in only one metric (accuracy or robustness). In this work, we examine nonlinear MBAs in general forms to show improvements in both metrics.
Analysis of Generalization
Demonstrating better generalization of classifiers often involves showing larger classification margins (Cao et al. 2019; Kawaguchi, Kaelbling, and Bengio 2017; Lim et al. 2021), but the proof for IPs is very different. A common but indirect approach shows an algorithm can regularize the Hessian of loss landscape with respect to the network parameters. Models that find flat minima in the loss landscape tend to generalize better in practice (Jiang et al. 2019; Keskar et al. 2016; Neyshabur et al. 2017; Foret et al. 2020). Many works demonstrate regularization of the Hessian (Xie, Sato, and Sugiyama 2020; Keskar et al. 2016; Foret et al. 2020) then complete the proof using the PAC-Bayes Theorem in hypothesis. While the PAC-Bayes theorem provides a bound on the generalization error, it does not directly incorporate curvature measures. The exact theoretical mechanism behind it is still not fully understood. In this work, instead of connecting to the PAC-Bayes Theorem, we rewrite the generalization risk and demonstrate that SGD jittering results in a smaller generalization risk than regular MSE training.
Noise Injections
Noise injection can be viewed as implicit regularization during training, which is employed to enhance model robustness and generalization. Input injection adds random noises to the signal space, which can be viewed as a data augmentation strategy. It is widely studied in classifications (Fawzi, Moosavi-Dezfooli, and Frossard 2016; Ford et al. 2019; Rusak et al. 2020), and recently shown to promote robustness in IPs (Krainovic, Soltanolkotabi, and Heckel 2023). Weight injection instead perturbed network parameters during training, which is shown to help escape from local minima (Zhu et al. 2018; Nguyen et al. 2019) and promote smooth solutions (Orvieto et al. 2023; Liu et al. 2021; Orvieto et al. 2022). Layer-wise noise injection (Lim et al. 2021) in a recurrent network has been shown to generalize better for classifiers. Weights can also be perturbed before parameter updates to enhance generalization, as in stochastic gradient descent (SGD). We will explore SGD in detail in the next subsection.
Connection to SGD Training
SGD training is effective in promoting generalization through implicit regularization (Smith et al. 2021). In large overparameterized models with non-unique global solutions, SGD favors networks with flat minima (Hochreiter and Schmidhuber 1997). The implicit regularization properties of SGD have been explored in contexts such as linear regression (Zou et al. 2021), logistic regression (Ji and Telgarsky 2018), and convolutional networks (Gunasekar et al. 2018). Unlike SGD training, which adds noise to network parameters, the proposed SGD jittering method introduces noise at each iterative update of the reconstructions in the lower-level optimization. Despite its name, SGD jittering is structurally similar to layer-wise noise injection within MBAs.
Definition of Robustness and Generalization
We formally define average-case robustness and generalization risks for a learned inverse mapping $H_{\theta}:{\bm{y}}\mapsto{\bm{x}}$ in IPs. Average-case robustness is also known as the “natural” robustness for average performance (Rice et al. 2021; Hendrycks and Dietterich 2019), which is equally important as in worst-case. During evaluation, a small vector ${\bm{g}}$ sampled from the distribution $\mathcal{P}_{g}$ is added to a data point ${\bm{x}}\in\mathcal{D}$. The new ground-truth becomes ${\bm{x}}_{g}={\bm{x}}+{\bm{g}}$, with the corresponding observation ${\bm{y}}_{g}=\bm{A}{\bm{x}}_{g}+{\bm{z}}={\bm{y}}+\bm{A}{\bm{g}}$.
Definition 1.
The average-case robustness risk of $H_{\theta}$ measures the distance between the unperturbed ${\bm{x}}$ to the recovered signal $\hat{{\bm{x}}}_{g}=H_{\theta}({\bm{y}}_{g})$ with the presence of ${\bm{g}}$,
$R_{e}(\theta)=\mathbb{E}_{{\bm{x}},{\bm{y}}\in\mathcal{D},{\bm{g}}\in\mathcal{P}_{g}}||{\bm{x}}-H_{\theta}({\bm{y}}_{g})||^{2}_{2}.$ (2)
Definition 2.
The generalization risk of $H_{\theta}$ measures the distance between the actual ground-truth ${\bm{x}}_{g}$ to the recovered signal $\hat{{\bm{x}}}_{g}=H_{\theta}({\bm{y}}_{g})$ with the presence of ${\bm{g}}$,
$\mathcal{G}(\theta)=\mathbb{E}_{{\bm{x}},{\bm{y}}\in\mathcal{D},{\bm{g}}\in\mathcal{P}_{g}}||{\bm{x}}_{g}-H_{\theta}({\bm{y}}_{g})||^{2}_{2}.$ (3)
Robustness measures how close the reconstruction is compared to the data within the dataset, and the generalization accuracy measures how well the reconstruction obeys the physics of the inverse problem even for out-of-distribution (OOD) data. Defining generalization in regression is not straightforward, but for IPs, we can achieve this using the underlying forward models. Typically, a tradeoff between robustness and accuracy exists for classical feedforward neural networks (Zhang et al. 2019; Yang et al. 2020; Stutz, Hein, and Schiele 2019). We aim to illustrate that the SGD jittering technique enhances robustness and accuracy for MBAs without incorporating explicit regularization.
Model-based Architecture Overview
MBAs are inspired by classical optimizations, where the underlying signal ${\bm{x}}$ can be estimated via,
$\hat{{\bm{x}}}=\operatorname*{arg\,min}_{\bm{x}}\frac{1}{2}||{\bm{y}}-\bm{A}{\bm{x}}||_{2}^{2}+r({\bm{x}}),$
where $r$ is an arbitrary regularization function to ensure the uniqueness of the reconstruction. When $r$ is differentiable, the optimal solution can be solved via gradient descent (GD), iterating through steps $k=1,2,3...$ with a constant step-size $\eta$,
${\bm{x}}_{k+1}={\bm{x}}_{k}-\eta\bm{A}^{\top}(\bm{A}{\bm{x}}_{k}-{\bm{y}})+\eta\nabla r({\bm{x}}_{k}).$
In a loop unrolling (LU) architecture using GD, $\nabla r$ is replaced by a shared-weight neural network, denoted by $f_{\theta}$. The process begins with an initial estimate ${\bm{x}}_{0}$, which is iteratively updated with a finite number of iterations. The final output ${\bm{x}}_{K}$ represents the predicted reconstruction and is compared to the desired target. Different variants of LU are summarized in Monga, Li, and Eldar (2021).
MBA Training Scheme Overview
In this section, we summarize some common training schemes for general regression problems, adapted here for MBA training. Following this, we introduce the proposed SGD jittering method for training MBAs.
Regular MSE Training
MSE training is the most commonly used training scheme for training regression problems and MBAs, where the goal using MBAs is to minimize the following risk w.r.t. $\theta$,
$R(\theta)$ $=\mathbb{E}_{({\bm{x}},{\bm{y}})\in\mathcal{D}}||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\quad\textrm{ where, }$, (4)
$\hat{{\bm{x}}}$ $=\operatorname*{arg\,min}_{x}~{}\frac{1}{2}\|{\bm{y}}-\bm{A}{\bm{x}}\|^{2}_{2}+r_{\theta}({\bm{x}})$
Notice that MBAs do to learn $r$ directly, but instead learn $\nabla r$ using a neural network $f_{\theta}$. We denote $r_{\theta}$ to highlight its dependency of $\theta$. The goal of MSE training is to minimize the squared distance between the output ${\bm{x}}_{K}$ and the ground-truth ${\bm{x}}$. MSE training might overfit the patterns in training data, making it more vulnerable to variations in ${\bm{y}}$.
Adversarial Training
AT is widely recognized as one of the most effective methods for training robust neural networks (Wang et al. 2021). AT finds the worst-case attack for each training instance and adds it to the input data designed to deceive the model. By learning from these challenging examples, AT improves the model’s robustness. AT aims to minimize the worst-case robustness risk for MBAs as follows,
$R_{\epsilon}(\theta)=\mathbb{E}_{({\bm{x}},{\bm{y}})\in\mathcal{D}}\left[\begin{aligned} &\max_{{\bm{e}}:||{\bm{e}}||_{2}\leq\epsilon}\quad||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\\ &\textrm{s.t. }\hat{{\bm{x}}}=\operatorname*{arg\,min}_{x}~{}\frac{1}{2}\|{\bm{y}}+{\bm{e}}-\bm{A}{\bm{x}}\|^{2}_{2}+r_{\theta}({\bm{x}})\end{aligned}\right].$ (5)
Fast gradient sign method (FGSM) and projected gradient descent (PGD) are methods to find the attack vector ${\bm{e}}$, where (Athalye, Carlini, and Wagner 2018) demonstrate that training a defense model with PGD adversarial attacks is more robust. However, AT is slow as it requires iterative solvers for the attack vector (Shafahi et al. 2019).
AT might suffer from bad generalizations. For example, in noiseless IPs, ${\bm{y}}=\bm{A}{\bm{x}}$, the true inverse $\bm{A}^{-1}$ exists. AT learns an inverse mapping $H_{\theta}$ so that $H_{\theta}({\bm{y}}+{\bm{e}})={\bm{x}}$, but the true estimated should be $\bm{A}^{-1}({\bm{y}}+{\bm{e}})\neq{\bm{x}}$.
Input Jittering
Noise injection to the input is widely used in general ML tasks to promote robustness (Salman et al. 2019; Cohen, Rosenfeld, and Kolter 2019; Lecuyer et al. 2019). In IPs that aim to map from ${\bm{y}}$ to ${\bm{x}}$, (Krainovic, Soltanolkotabi, and Heckel 2023) proposes to add zero-mean random Gaussian vector ${\bm{w}}$ in ${\bm{y}}$ to a neural network that learns a direct reconstruction, and demonstrate that with careful selection of the jittering variance, this training scheme can attain a comparable level of worst-case robustness to that achieved by AT. In MBAs, the input jittering approach is expressed as,
$J^{I}_{\sigma_{w}}(\theta)$ $=\mathbb{E}_{{\bm{w}},({\bm{x}},{\bm{y}})\in\mathcal{D}}||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\quad\textrm{ where,}$, (6)
$\hat{{\bm{x}}}$ $=\operatorname*{arg\,min}_{x}~{}\frac{1}{2}\|{\bm{y}}+{\bm{w}}-\bm{A}{\bm{x}}\|^{2}_{2}+r_{\theta}({\bm{x}}).$
This method learns an inverse mapping from ${\bm{y}}+{\bm{w}}$ to ${\bm{x}}$, which might suffer from generalization issues in the same way as AT. In fact, Krainovic, Soltanolkotabi, and Heckel (2023) demonstrates a significant accuracy loss using input jittering for black-box neural networks.
Proposed Training Scheme
Another line of work injects random noises layer-wise into the neural network, where (Hodgkinson et al. 2021; Jim, Giles, and Horne 1996; Liu et al. 2020) model the layer-wise injection as stochastic differential equation and show implicit regularization in robustness, and (Lim et al. 2021) added noises to the recurrent neural network and demonstrates better generalization for classification tasks.
In this work, we propose adding random noise vectors ${\bm{w}}_{k}$ drawn from a zero-mean Gaussian distribution to the learned gradient updates $f_{\theta}({\bm{x}}_{k})$, with the noise being re-sampled independently for each iteration $k=1,2,3,...$. The noisy gradients are unbiased, $\mathbb{E}[f_{\theta}({\bm{x}})+{\bm{w}}_{k}]=\mathbb{E}[f_{\theta}({\bm{x}})]$ for all ${\bm{x}}$. Let $\bar{{\bm{w}}}=\{{\bm{w}}_{1},...,{\bm{w}}_{K}\}$. The estimated reconstruction is solved iteratively using this noisy gradient descent which is essentially the stochastic gradient descent (SGD) with step size $\eta$. Thus, the learning objective is as follows,
$J^{SGD}_{\sigma_{w_{k}}}(\theta)$ $=\mathbb{E}_{\bar{{\bm{w}}},({\bm{x}},{\bm{y}})\in\mathcal{D}}||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\quad\textrm{ where,}$, (7)
$\hat{{\bm{x}}}$ $=\lim_{k\rightarrow\infty}{\bm{x}}_{k}-\eta\big{(}\bm{A}^{\top}(\bm{A}{\bm{x}}_{k}-{\bm{y}})+f_{\theta}({\bm{x}}_{k})+{\bm{w}}_{k}\big{)}$
In AT and input jittering training, the noise vectors are added to ${\bm{y}}$ throughout the entire inversion, introducing a bias to the reconstruction compared to the actual ground-truth. In contrast, SGD jittering adds noise to the learned gradient descent updates. The lower-level problem in SGD jittering minimizes $\frac{1}{2}||{\bm{y}}-\bm{A}{\bm{x}}||_{2}^{2}+r_{\theta}({\bm{x}})$, with the convergence result in Lemma 1 ensuring the correct inverse mapping even with noise in gradient descent.
We modified the convergence result of the SGD in (Ghadimi and Lan 2013) for lower-level SGD in MBAs as follows, where the proof of Lemma 1 can be found in the Appendix.
Assumption 1.
(SGD noises) Assume noises ${\bm{w}}_{k}\sim\mathcal{N}(0,\sigma_{w_{k}}^{2}/n\mathbf{I})$ sampled from i.i.d. zero-mean Gaussian distribution for each $k=1,2,3,...$ are added to each gradient descent iteration, so that $\mathbb{E}||{\bm{w}}_{k}||_{2}^{2}=\sigma_{w_{k}}^{2}$.
Lemma 1.
(Convergence of MBAs-SGD for IPs, modified from SGD convergence result in (Ghadimi and Lan 2013)) Let $F({\bm{x}})=\frac{1}{2}||{\bm{y}}-\bm{A}{\bm{x}}||_{2}^{2}+r({\bm{x}})$ denote the lower-level objective function. Assume $f_{\theta}({\bm{x}})$ is $L$-Lipschitz continuous,
$||f_{\theta}({\bm{x}}+\Delta)-f_{\theta}({\bm{x}})||_{2}\leq L||\Delta||_{2},\quad\forall{\bm{x}},\Delta$
With assumption 1, and step-sizes $\eta_{k}<2/L$, SGD in minimizing $F({\bm{x}})$ converges to a stationary point.
SGD noise also implicitly enhances robustness. Intuitively, exposing a neural network to noisier inputs during training encourages it to converge to smoother regions of the loss landscape, which are less sensitive to input changes. This leads to more stable outputs that are less affected by perturbations during evaluation. We will analyze how SGD jittering contributes to this robustness in the following section.
Analysis of Denoising Problem
In this section, we analyze the generalization and robustness of SGD jittering in addressing the denoising problem. We begin by outlining the relevant assumptions.
Assumption 2.
(Denoising) The forward model in denoising is identity mapping, so that ${\bm{y}}={\bm{x}}+{\bm{z}}$. Signals ${\bm{x}}\in\mathbb{R}^{n}$ are sampled from distribution $\mathcal{X}$, and noises are sampled i.i.d. from ${\bm{z}}\sim\mathcal{N}(0,\sigma_{z}^{2}/n\bm{I})$.
Assumption 3.
(Neural Network) Assume $f_{\theta}$ is twice differentiable with respect to the input, so the Hessian exists.
Figure 1: 2D point denoising example using MSE training, AT, and SGD jittering. Each method is evaluated using regular in-distribution data, under average- and worst-case attacks to ${\bm{y}}$ and with additional bias (from left to right columns respectively).
Main Theoretical Results
Given the assumptions, we state the first theorem regarding generalization accuracy. Let $H_{\theta}$ denote the inverse process learned using MBAs for the rest of the analysis.
Theorem 2.
Assuming a small zero-meam Gaussian random vector ${\bm{g}}$ with $\mathbb{E}||{\bm{g}}||^{2}_{2}=\sigma_{g}^{2}$ is added to data ${\bm{x}}\in\mathcal{D}$ during evaluation. The generalization risk for denoising problem is defined as,
$\mathcal{G}(\theta)=\mathbb{E}_{{\bm{x}},{\bm{y}}\in\mathcal{D},{\bm{g}}}||{\bm{x}}+{\bm{g}}-H_{\theta}({\bm{y}}+{\bm{g}})||^{2}_{2}.$ (8)
Under assumptions 1, 2 and 3, SGD jittering generalizes better than regular MSE training,
$\mathcal{G}(\theta^{sgd})\leq\mathcal{G}(\theta^{mse}).$
Proof.
For any trained model $\theta$, we have generalization risk in (8). We can express $H_{\theta}({\bm{y}}+{\bm{g}})$ in terms of $H_{\theta}({\bm{y}})$ for the same set of parameters via iterative expansions. For iterations $k=0,1,...,K$, let the sequence $\{{\bm{x}}_{0},{\bm{x}}_{1},...,{\bm{x}}_{K}\}$ denote the denoising trajectory from observation ${\bm{y}}+{\bm{g}}$, where $H_{\theta}({\bm{y}}+{\bm{g}})={\bm{x}}_{K}$, and let $\{{\bm{x}}_{0}^{\prime},{\bm{x}}_{1}^{\prime},...,{\bm{x}}_{K}^{\prime}\}$ denote the denoising trajectory from observation ${\bm{y}}$ where $H_{\theta}({\bm{y}})={\bm{x}}_{K}^{\prime}$. Then we can rewrite the reconstruction trajectory with the appearance of ${\bm{g}}$ as follows,
${\bm{x}}_{0}={\bm{x}}_{0}^{\prime}+{\bm{g}}={\bm{y}}+{\bm{g}}$
${\bm{x}}_{1}=(1-\eta){\bm{x}}_{0}+\eta({\bm{y}}+{\bm{g}})-\eta f_{\theta}({\bm{x}}_{0})$
$=\big{(}(1-\eta){\bm{x}}_{0}^{\prime}+\eta{\bm{y}}-\eta f_{\theta}({\bm{x}}_{0}^{\prime})\big{)}+\big{(}{\bm{g}}+\eta f_{\theta}({\bm{x}}_{0}^{\prime})-\eta f_{\theta}({\bm{x}}_{0})\big{)}$
$={\bm{x}}_{1}^{\prime}+{\bm{g}}+\eta(f_{\theta}({\bm{x}}_{0}^{\prime})-f_{\theta}({\bm{x}}_{0}))$
${\bm{x}}_{2}=(1-\eta){\bm{x}}_{1}+\eta({\bm{y}}+{\bm{g}})-\eta f_{\theta}({\bm{x}}_{1})$
$={\bm{x}}_{2}^{\prime}+{\bm{g}}+\eta(1-\eta)(f_{\theta}({\bm{x}}_{0}^{\prime})-f_{\theta}({\bm{x}}_{0}))+\eta(f_{\theta}({\bm{x}}_{1}^{\prime})-f_{\theta}({\bm{x}}_{1}))$
$\quad$...
${\bm{x}}_{k+1}={\bm{x}}_{k+1}^{\prime}+{\bm{g}}+\sum_{i=0}^{k}\eta(1-\eta)^{k-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}))$
Let $K=k+1$ and plug in the above expansion to (8). For brevity, we omit the subscripts of expectations in the proof.
$\mathcal{G}(\theta)=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}-\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}))||^{2}_{2}$
$=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}))||^{2}_{2}$
The first term is the testing accuracy when ${\bm{x}},{\bm{y}}$ are in the data distribution, while the second term is affected by ${\bm{g}}$. MSE training minimizes the first term solely, but no extra regularization on the second term. On the other hand, we will show that SGD jittering training also penalizes the magnitude of the second term, thus obtaining a smaller generalization risk than MSE training.
We view the SGD jittering process as a noisy version of regular MSE training, the goal is to write the SGD jittering risk in terms of the noiseless updates with some implicit regularization. Let $\{{\bm{x}}_{0}^{\prime},{\bm{x}}_{1}^{\prime},...,{\bm{x}}_{K}^{\prime}\}$ denote the noiseless GD trajectory of a model $\theta$. We write the iterative updates of SGD jittering training $\{{\bm{x}}_{0}^{sgd},{\bm{x}}_{1}^{sgd},...,{\bm{x}}_{K}^{sgd}\}$ for the same set of parameters $\theta$ in terms of ${\bm{x}}_{k}^{\prime}$s, which tells the deviation of each intermediate reconstruction between the SGD jittering and its noiseless counterpart,
${\bm{x}}_{0}^{sgd}={\bm{y}}={\bm{x}}_{0}^{\prime}+{\bm{w}}_{0}$
${\bm{x}}_{1}^{sgd}=(1-\eta){\bm{x}}_{0}^{sgd}+\eta{\bm{y}}-\eta f_{\theta}({\bm{x}}_{0}^{sgd})-\eta{\bm{w}}_{1}$
$={\bm{x}}_{1}^{\prime}+\eta(f_{\theta}({\bm{x}}_{0}^{\prime})-f_{\theta}({\bm{x}}_{0}^{sgd}))-(\eta(1-\eta){\bm{w}}_{0}+\eta{\bm{w}}_{1})$
${\bm{x}}_{2}^{sgd}=(1-\eta){\bm{x}}_{1}^{sgd}+\eta{\bm{y}}-\eta f_{\theta}({\bm{x}}_{1}^{sgd})-\eta{\bm{w}}_{2}$
$={\bm{x}}_{2}^{\prime}+\eta\big{(}f_{\theta}({\bm{x}}_{1}^{\prime})-f_{\theta}({\bm{x}}_{1}^{sgd})\big{)}+\eta(1-\eta)\big{(}f_{\theta}({\bm{x}}_{0}^{\prime})-f_{\theta}({\bm{x}}_{0}^{sgd})\big{)}$
$\qquad-\big{(}\eta(1-\eta)^{2}{\bm{w}}_{0}+\eta(1-\eta){\bm{w}}_{1}+\eta{\bm{w}}_{2}\big{)}$
${\bm{x}}_{k+1}^{sgd}=(1-\eta){\bm{x}}_{k}^{sgd}+\eta{\bm{y}}-\eta f_{\theta}({\bm{x}}_{k}^{sgd})-\eta{\bm{w}}_{k+1}$
$={\bm{x}}_{k+1}^{\prime}+\sum_{i=0}^{k}\eta(1-\eta)^{k-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd}))$
$\qquad-\sum_{i=0}^{k+1}\eta(1-\eta)^{k+1-i}{\bm{w}}_{i}.$
Let $K=k+1$, SGD jittering training risk is,
$\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{sgd}||_{2}^{2}=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}~{}||\sum_{i=0}^{K}\eta(1-\eta)^{K-i}{\bm{w}}_{i}||_{2}^{2}$
$+\mathbb{E}~{}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd}))||_{2}^{2}$
$-2\mathbb{E}\left[\langle\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}f_{\theta}({\bm{x}}_{i}^{sgd}),\sum_{i=0}^{K}\eta(1-\eta)^{K-i}{\bm{w}}_{i}\rangle\right].$
The first component corresponds to the MSE loss when the trajectory is not perturbed by jittering noise. The second term is weighted noise variance independent of the network parameters. The third term adds extra regularization to the difference in outputs of $f_{\theta}$ between noisy and clean trajectories. It penalizes more as the iterations approach the final output. The last term is the cross product between the jittering noise and the perturbed reconstructions from $f$, which is lower bounded since the overall SGD jittering risk is non-negative.
When choosing $\sigma_{{\bm{w}}_{k}}^{2}$ at iteration $k$ such that $\sum_{i=0}^{k}\eta^{2}(1-\eta)^{2(k-i)}\sigma_{w_{i}}^{2}=\sigma_{\bm{g}}^{2}$, the reconstruction from ${\bm{x}}+{\bm{g}}$ is a special case in SGD jittering training where ${\bm{w}}_{k}$ are independently sampled at each iteration. Thus, minimizing the regularization term $\mathbb{E}_{({\bm{x}},{\bm{y}})\sim\mathcal{D},{\bm{w}}_{k}\forall k}~{}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd}))||_{2}^{2}$ also reduces the penalty term in computing generalization accuracy $\mathbb{E}_{{\bm{x}},{\bm{y}}\in\mathcal{D},{\bm{g}}}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}))||^{2}_{2}$. Therefore, $\mathcal{G}(\theta^{sgd})\leq\mathcal{G}(\theta^{mse})$ due to the implicit regularization term in SGD jittering training.
In addition, we can expand the regularization term in SGD training via Taylor expansion. Let $\bm{\delta}_{k+1}=\sum_{i=0}^{k}\eta(1-\eta)^{k-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd}))+\sum_{i=0}^{k+1}\eta(1-\eta)^{k+1-i}{\bm{w}}_{i}$. Assumption 3 ensures the existence of the gradient and the Hessian of $f_{\theta}$. We have,
$f_{\theta}({\bm{x}}_{k+1}^{sgd})\approx f_{\theta}({\bm{x}}_{k+1}^{\prime})+\bm{\delta}_{k+1}^{\top}\nabla_{\bm{x}}f_{\theta}({\bm{x}}_{k+1}^{\prime})+\frac{1}{2}\bm{\delta}_{k+1}^{\top}\nabla^{2}_{\bm{x}}f_{\theta}({\bm{x}}_{k+1}^{\prime})\bm{\delta}_{k+1}.$
The higher the iteration $k$, the larger penalty weight $\eta(1-\eta)^{K-1-i}$ becomes. Thus with a large $K$, SGD jittering promotes a smoother and wider landscape for the iterative inputs of the $f_{\theta}$. We also refer to this as a flat minimum because the update of ${\bm{x}}_{k}$ solves the lower-level minimization problem. This is consistent with the observations in (Lim et al. 2021), which demonstrate that adding random Gaussian noise to recurrent neural networks layer-wise promotes a small Hessian in the last single-block neural network in terms of its inputs. While it is widely observed that flat minima in the loss landscape (with respect to network parameters) imply better generalization as summarized in Related Works, (Lim et al. 2021) does not clarify why flat minima with respect to hidden-layer inputs improve generalization in regression settings. Our work fills this gap in the context of solving IPs. Now, we state the robustness result,
Theorem 3.
Let ${\bm{e}}$ be the perturbation vector sampled i.i.d. from the zero-mean Gaussian distribution, $\mathcal{P}_{e}$, such that $\mathbb{E}||{\bm{e}}||_{2}^{2}=\sigma_{\bm{e}}^{2}$. For an inverse mapping $H_{\theta}$, the average-case robustness risk for denoising is as follows,
$R_{e}(\theta)=\mathbb{E}_{({\bm{x}},{\bm{y}})\in\mathcal{D},{\bm{e}}\in\mathcal{P}_{e}}~{}\left[||{\bm{x}}-H_{\theta}({\bm{y}}+{\bm{e}})||^{2}_{2}\right].$ (9)
SGD jittering is more robust than MSE training against bounded-variance perturbations around the measurements,
$R_{e}(\theta^{sgd})\leq R_{e}(\theta^{mse}).$
The proof is similar to the one in proving generalization, we include the full proof in the Appendix. The gradient and Hessian of $f_{\theta}$ together help characterize the generalization and robustness of MBAs.
Figure 2: Seismic deconvolution under adversarial attack.
Experiments and Discussion
In our experiments, we use a toy denoising example to validate our theorems and demonstrate the proposed training scheme using seismic deconvolution and magnetic resonance imaging (MRI) reconstruction. For each task, we train the same LU architecture using different training schemes. We then evaluate the models using in-distribution testing data, adversarial attacks with various strengths, and task-dependent OOD data. Detailed training procedures are provided in the Appendix.
Toy problem: 2-dimensional Point Denoising
To validate the Theorems, we compare the robustness and generalization using regular MSE, AT and SGD-jittering in a point denoising problem. Assume ${\bm{x}}=(0,0)^{\top}$, we generate noisy ${\bm{y}}$ by adding random zero-mean Gaussian noise with a variance of 0.01. We use 200 samples for training and 50 for evaluation. Figure 1 shows the noisy points ${\bm{y}}$ and the denoised results $\hat{{\bm{x}}}$ (blue) across four evaluation modes, with the ground-truth in red. In-distribution data are sampled from the same distribution as the training data, centered at $(0,0)^{\top}$. The worst-case attack finds a vector within a 0.01 $\ell_{2}$ distance around ${\bm{y}}$ that maximizes the squared error, while the average-case attack uniformly samples attack vectors within the same ball and adds to ${\bm{y}}$. SGD jittering and AT are more robust than MSE training in both metrics. AT aims to promote worst-case robustness, which implies average-case robustness for attacks within the same $\ell_{2}$ ball around ${\bm{y}}$, though its average-case performance may not be optimal. We observe that SGD jittering has even tighter clusters in its reconstructions under average-case attacks than AT. Finally, we evaluate generalization accuracy by reconstructing the slightly biased ground truth $(0.005,0)^{\top}$ from its corresponding measurements. With the introduction of bias, SGD jittering achieves a more accurate reconstruction compared to the new ground truth. Notice that MSE training underperforms on in-distribution data compared to SGD jittering, likely because the noise injection in SGD jittering helps the model escape from local minima in training.
Figure 3: Seismic deconvolution with an additional layer.
Figure 4: MRI evaluation (a) under adversarial attack obtained from projected gradient descent with norm no more than $\epsilon=1$. Each row represents an MRI reconstruction of a sample from fastMRI (Zbontar et al. 2018a). (b) MRI evaluation using OOD tumor knee data from (Bickle and Jin 2021).
Seismic Deconvolution
Seismic deconvolution is a crucial problem in geophysics. It aims to reverse the convolution effects in the received signal, which occurs when artificial source waves travel through the Earth’s layers. The goal is to extract sparse reflectivity series from recorded seismic data. The training data is generated following the same procedure as in (Iqbal et al. 2019; Guan et al. 2024). To demonstrate the generalization accuracy, we generate OOD testing data by adding an arbitrary horizontal layer with magnitudes of 0.1 and 0.3 to the ground truth, reflected accordingly in the measurements. Table 1 numerically compares the average testing peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) of the reconstructions. Notice that for OOD data, the numerical values are very similar, as the additional layer has a small magnitude relative to the overall measurement. However, the difference can be clearly observed in the figures. Figure 2 shows two deconvolution examples for each trained model under adversarial attacks. MSE training produces artifacts, while both AT and SGD jittering yield results closer to the ground truth. Notably, in the top example, AT fails to recover the small horizontal layer in the center of the boxed area, but SGD jittering successfully captures it. Figure 3 demonstrates the generalization ability with two different magnitude levels. For the OOD example with the larger magnitude, both SGD jittering and MSE training recover a cleaner layer compared to AT. More notably, with the smaller magnitude, AT perceives the line as an unwanted perturbation and barely recovers it, whereas the result from SGD jittering is more distinctly visible.
PSNR / SSIM Test data Adv. Attack OOD small OOD large
MSE training 34.64 / 0.921 28.85 / 0.829 34.57 / 0.918 34.57 / 0.917
AT 33.50 / 0.903 30.27 / 0.849 33.45 / 0.902 33.43 / 0.902
Input Jittering 32.92 / 0.882 30.10 / 0.832 32.91 / 0.882 32.90 / 0.882
SGD jittering 35.10 / 0.928 29.89 / 0.842 34.93 / 0.927 34.93 / 0.927
Table 1: Seismic deconvolution evaluation for in-distribution data (column 2), under adversarial attack with $\epsilon=1$ (column 3), and for OOD data with small and large magnitudes (columns 4 and 5). The best and second-best performances are in bold and underlined respectively.
Accelerated MRI Reconstruction
Accelerated MRI aims to recover human-interpreted body structures from partial k-space measurements. We train models with single-coil knee MRI from the fastMRI dataset (Zbontar et al. 2018b) with 4×4$\times$4 × acceleration, or $1/4$ of the measurements in k-space is used for reconstruction. To assess generalization, we use a different knee dataset from Bickle and Jin (2021), which includes giant tumor cells not present in the training dataset. Table 2 numerically compares the average testing PSNR and SSIM of the reconstructions for in-distribution fastMRI data, under adversarial attack with $\epsilon=1$ and OOD tumor cell knee data. Our findings indicate that MSE training and SGD jittering exhibit comparable performance on in-distribution data, but SGD jittering demonstrates greater robustness and superior performance on OOD data. Adversarial training (AT) yields the most robust model under attack, though it comes with a significant tradeoff in performance for both in-distribution and OOD data. Additionally, as shown in Figure 4, reconstructions using AT appear blurrier than others, a similar observation noted in black-box IP solvers by Krainovic, Soltanolkotabi, and Heckel (2023).
PSNR / SSIM fastMRI Adv. Attack Tumor Cell
MSE training 28.21 / 0.603 25.68 / 0.382 29.92 / 0.779
AT 27.68 / 0.564 27.17 / 0.549 27.74 / 0.597
Input Jittering 28.18 / 0.595 25.05 / 0.420 29.97 / 0.740
SGD jittering 28.22 / 0.607 26.77 / 0.552 30.36 / 0.788
Table 2: MRI evaluation for in-distribution data (column 2), adversarial attack (column 3), and OOD data (column 4). The best and second best performances are in bold and underlined respectively.
Conclusion
Robustness and generalization accuracy are both crucial in solving IPs, yet current analyses often focus on simple black-box solvers that enhance robustness at the cost of accuracy. While empirical studies have examined the robustness of MBAs, there is a lack of theoretical analysis. In this work, we introduce a novel and easily implemented SGD jittering training scheme to address the tradeoff between robustness and accuracy in solving IPs. Our method proves effective for high-quality model-based IP solvers. Through mathematical analysis, we demonstrate that this approach implicitly regularizes the gradient and Hessian of the neural network with respect to the input, resulting in improved generalization and enhanced robustness compared to traditional MSE training. Experimental results show that while MSE training is vulnerable to adversarial attacks and AT recovers smooth estimates, SGD jittering consistently produces robust, high-quality outcomes.
Appendix A Appendix
Appendix B Notations
We clarify the notations used in the paper.
  • $H_{\theta}$: learned inverse mapping parameterized by $\theta$, i.e., black-box neural networks, MBAs. For MBAs in particular, $H_{\theta}=\operatorname*{arg\,min}_{x}~{}\frac{1}{2}\|{\bm{y}}-\bm{A}{\bm{x}}\|^{2}_{2}+r_{\theta}({\bm{x}})$.
  • $r$: regularization function from predefined functions, i.e., $\ell_{1}$ and $\ell_{2}$ norm.
  • $r_{\theta}$: regularization function whose gradient is learned from a neural network $f_{\theta}$.
  • $f_{\theta}$: learned gradient update of $r$, $f_{\theta}({\bm{x}})=\nabla_{\bm{x}}r_{\theta}({\bm{x}})$.
  • $F(x)=\frac{1}{2}\|{\bm{y}}-\bm{A}{\bm{x}}\|^{2}_{2}+r_{\theta}({\bm{x}})$, denote the value of the lower-level objective function.
Appendix C Proof of Lemma 1
Lemma 4.
(Restate Lemma 1. Convergence of MBAs-SGD for IPs) Let $F({\bm{x}})=\frac{1}{2}||{\bm{y}}-\bm{A}{\bm{x}}||_{2}^{2}+r({\bm{x}})$ denote the lower-level objective function. Assume $f_{\theta}({\bm{x}})$ is $L$-Lipschitz continuous,
$||f_{\theta}({\bm{x}}+\Delta)-f_{\theta}({\bm{x}})||_{2}\leq L||\Delta||_{2},\quad\forall{\bm{x}},\Delta$
With assumption 1, and step-sizes $\eta_{k}<2/L$, SGD in minimizing $F({\bm{x}})$ converges to a stationary point.
Proof.
The lemma can be easily derived from the Theorem in (Ghadimi and Lan 2013), which assumes $\nabla_{\bm{x}}F({\bm{x}})$ is $L$-Lipschitz continuous.
$||\nabla_{\bm{x}}F({\bm{x}}+\Delta)-\nabla_{\bm{x}}F({\bm{x}})||_{2}$
$=$ $||\bm{A}^{\top}(\bm{A}({\bm{x}}+\Delta)-{\bm{y}})+f_{\theta}({\bm{x}}+\Delta)-\bm{A}^{\top}(\bm{A}{\bm{x}}-{\bm{y}})-f_{\theta}({\bm{x}})||_{2}$
$=$ $||\bm{A}^{\top}\bm{A}\Delta+f_{\theta}({\bm{x}}+\Delta)-f_{\theta}({\bm{x}})||_{2}$
$\leq$ $||\bm{A}^{\top}\bm{A}||\,||\Delta||_{2}+||f_{\theta}({\bm{x}}+\Delta)-f_{\theta}({\bm{x}})||_{2}$
$\leq$ $\max_{i=1,..,n}\lambda_{i}||\Delta||_{2}+||f_{\theta}({\bm{x}}+\Delta)-f_{\theta}({\bm{x}})||_{2}$
where $\lambda_{i}$ be the eigenvalues of $\bm{A}^{\top}\bm{A}$, and let $\mu=\max_{i=1,..,n}\lambda_{i}$. Since $\nabla F$ is $L$-Lipshitz, we show that for MBAs-SGD in solving denoising problems, it is equivalent to $(L-\mu)$-Lipschitz continuous of $f_{\theta}$. With the same assumptions on step sizes and the SGD jittering noise, we finish the proof. ∎
Appendix D Proof of Theorem 3 - robustness
Theorem 5.
(Restate Theorem 3) Let ${\bm{e}}$ be the perturbation vector and let $\mathcal{P}_{e}$ be iid zero-mean Gaussian distribution such that $\mathbb{E}||{\bm{e}}||_{2}^{2}=\sigma_{\bm{e}}^{2}$. The average-case robustness risk for denoising problem is as follows,
$R_{e}(\theta)=\mathbb{E}_{({\bm{x}},{\bm{y}})\in\mathcal{D},{\bm{e}}}~{}\left[||{\bm{x}}-H_{\theta}({\bm{y}}+{\bm{e}})||^{2}_{2}\right].$ (10)
For the denoising problem, SGD jittering is more robust than MSE training against bounded-variance perturbations around the measurement, or
$R_{e}(\theta^{sgd})\leq R_{e}(\theta^{mse}).$
Proof.
The proof is similar to proving generalization accuracy. To evaluate the average-case robustness for any trained $\theta$, we first write the iterative updates under attack $\{{\bm{x}}_{0}^{at},{\bm{x}}_{1}^{at},...{\bm{x}}_{K}^{at}\}$ in terms of the noiseless trajectory $\{{\bm{x}}_{0}^{\prime},{\bm{x}}_{1}^{\prime},...{\bm{x}}_{K}^{\prime}\}$.
${\bm{x}}_{0}^{at}$ $={\bm{x}}_{0}^{\prime}+{\bm{e}}={\bm{y}}+{\bm{e}}$
${\bm{x}}_{1}^{at}$ $=(1-\eta){\bm{x}}_{0}^{at}+\eta({\bm{y}}+{\bm{e}})-\eta f_{\theta}({\bm{x}}_{0}^{at})$
$=$ ${\bm{x}}_{1}^{\prime}+{\bm{e}}+\eta(f_{\theta}({\bm{x}}_{0}^{\prime})-f_{\theta}({\bm{x}}_{0}^{at}))$
${\bm{x}}_{2}^{at}$ $=(1-\eta){\bm{x}}_{1}^{at}+\eta({\bm{y}}+{\bm{e}})-\eta f_{\theta}({\bm{x}}_{1}^{at})$
$=$ ${\bm{x}}_{2}^{\prime}+{\bm{e}}+\eta(1-\eta)(f_{\theta}({\bm{x}}_{0}^{\prime})-f_{\theta}({\bm{x}}_{0}^{at}))+\eta(f_{\theta}({\bm{x}}_{1}^{\prime})-f_{\theta}({\bm{x}}_{1}^{at}))$
$\quad$...
${\bm{x}}_{k+1}^{at}$ $={\bm{x}}_{k+1}^{\prime}+{\bm{e}}+\sum_{i=0}^{k}\eta(1-\eta)^{k-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))$
Let $K=k+1$, we rewrite the robustness risk. Notice that the generalization loss in (8) includes the additional vector ${\bm{g}}$, whereas the robustness risk in (9) does not.
$R_{e}(\theta)=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}-{\bm{e}}-\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))||^{2}_{2}$
$=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))||^{2}_{2}$
$\quad+\mathbb{E}||{\bm{e}}||_{2}^{2}+2\mathbb{E}\left[<{\bm{e}},\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))>\right]$
$=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))||^{2}_{2}$
$\quad+\mathbb{E}||{\bm{e}}||_{2}^{2}+2\mathbb{E}\left[<{\bm{e}},-\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}-f_{\theta}({\bm{x}}_{i}^{at})>\right]$
$=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))||^{2}_{2}+\sigma_{\bm{e}}^{2}$
$\quad+2\mathbb{E}\left[\langle{\bm{e}},-\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}-f_{\theta}({\bm{x}}_{i}^{at})\rangle\right]$
The first term is the mean-squared testing loss for in-distribution data. The second term is an additional penalty that appears in SGD jittering training. When picking $\sigma_{{\bm{w}}_{k}}^{2}$ for all $k$ and $\sigma_{\bm{e}}^{2}$ such that $\sum_{k=0}^{k}\eta^{2}(1-\eta)^{2(k-i)}\sigma_{w_{i}}^{2}=\sigma_{\bm{e}}^{2}$, this term in robustness risk is a special case in SGD jittering risk. The last term also appears in SGD jittering risk, when the above condition for jittering level and robustness level are satisfied, ${\bm{e}}$ and $\sum_{i=0}^{K}\eta(1-\eta)^{K-i}{\bm{w}}_{i}$ are from the same distribution. Therefore, for models that perform equally well in mean-squared error for in-distribution testing data, minimizing SGD jittering risk minimizes overall robustness. ∎
Appendix E Training Details
We use 10-iteration gradient descent loop unrolling architectures for all tasks. For the toy example, we use a 3-layer MLP with a hidden dimension of 32 as the learned gradient network. For seismic deconvolution and MRI reconstruction, a 5-layer and 8-layer DnCNN with 64 hidden channels are used for the learned gradient network, respectively. All models are trained using Nvidia RTX3080, using Adam optimizer with a learning rate of $41e-$ 4.
The jittering noise variance for each method is selected based on performances of both robustness and accuracy. In input jittering training, the three methods in order of 2D denoising, seismic deconvolution and MRI reconstruction use the noise variance of 0.01, 0.05 and 0.05. In SGD jittering training, we choose the DGD jittering variance for each task with 0.01, 0.1 and 0.01 respectively.
Appendix F Training Speed
We also compare the training speed among different training schemes. All methods are trained on the same device. Training batch sizes for the 2D denoising problem, seismic deconvolution and MRI are 256, 16, and 4 respectively. AT is trained by projected gradient descent. For the 2D denoising problem, $\epsilon=0.01$ with 50 PGD iterations and a stepsize of 0.05. For the seismic deconvolution and MRI reconstruction problem, $\epsilon=1$, with 20 PGD iterations and a stepsize of 0.1.
2D denoising Seis. Deconv. MRI
MSE training 5360.84 it/s 23.69 it/s 5.23 it/s
AT 562.75 it/s 1.53 it/s 0.25 it/s
Input Jittering 5256.77 it/s 22.80 it/s 5.20 it/s
SGD jittering 4337.84 it/s 22.72 it/s 5.27 it/s
Table 3: Training speed measured in items per second are recorded.
AT is significantly slower than other methods. Notice that SGD jittering is very time efficient, even with the additional noise sampling steps.